OpenStudy (anonymous):

what is 2^x=10 solve for x please show steps

OpenStudy (regan):

2=e^ln(2) e^ln(2)x=10 ln(e^ln(2)x)=ln(10) ln(2)x=ln(10) x=ln(10)/ln(2)

OpenStudy (anonymous):

can you explain it please

OpenStudy (regan):

do you know what a logarithm is?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

is that what ln stands foor

OpenStudy (regan):

kind of. ok imagine if we guess some values for x that will be close to 10 2^2=4 2^3=8 2^4=16 so we can imagine x lies in between 3 and 4 do you get that?

OpenStudy (anonymous):

yes

OpenStudy (regan):

so we need a method of finding x accurately. so we use logs. logs shift us between bases and powers. we 'log' both sides just as we would when using any function

OpenStudy (regan):

log(2^x)=log(10) the law of exponents say log(2^x)=xlog(2)

OpenStudy (regan):

log(10)/log(2)=x try it out on a calculator the answer will be inbetween 3 and 4